Which of the following are the first three iterates of the function f(x)=3x+2 for the initial value of x^0=-2
a. -2,-4,-10
b. 4, -10,28
c. -4, -10, -28
d. -4, 10, -28
I think the answer is (C). Is that correct?

Answer by jim_thompson5910(35256)

I'm not 100% familiar with the topic of function iterates, but my guess is that you plug in an initial value x = x0 to get x1 = f(x0). Then you plug in x1 = f(x0) to get x2 = f(f(x0)), etc etc to generate a sequence of values.


If that's the case, then you are correct.